# On the number of modes of Gaussian kernel density estimators

### v2

@@ 1.4 Notation
We adopt standard notation from asymptotic analysis: we write $f(x)\ll g(x)$ or $f(x)=o(g(x))$ if $f(x)/g(x)\to 0$ as $x\to\infty$; $f(x)\lesssim g(x)$ or $f(x)=O(g(x))$ if there exists a finite, positive constant $C$ such that $f(x)\leqslant Cg(x)$; and we write $f(x)\asymp g(x)$ or $f(x)=\Theta(g(x))$ if $f(x)\lesssim g(x)$ and $g(x)\lesssim f(x)$.
@@ 2.1 The Kac-Rice formula
Theorem 2.1 (Kac-Rice, [AW09, pp. 62], [AT09, Section 11.1]).
@@
    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity111defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$. of $\Psi^{\prime}(\cdot)$.
@@
The Kac-Rice formula extends to any dimension $d\geqslant 1$, and also on manifolds other than $\mathbb{R}^{d}$—see [AT09, Section 11.1]. It is the classical tool for computing the expected number of critical points of random fields, with many recent applications including spin glasses [AAČ13, FMM21] and landscapes of loss functions arising in machine learning [MAB20]. While the method applies to general densities, the conditional expectation appears infeasible to compute or estimate beyond the Gaussian case.
For the KDE $\widehat{P}_{n}$ defined in 1.1, define the random function $F_{n}:\mathbb{R}\to\mathbb{R}$ by
@@
Then $t\in\mathbb{R}$ is an upcrossing of $F_{n}$ at level 00 if and only if $F_{n}(t)=0$ and $F^{\prime}_{n}(t)>0$. This is equivalent to $\widehat{P}^{\prime}_{n}(t)=0$ and $\widehat{P}_{n}^{\prime\prime}(t)<0$, i.e. $t$ is a mode of $\widehat{P}_{n}$. Thus, the number of modes of $\widehat{P}_{n}$ in $T$ is given by $U_{0}(F_{n},T)$. For $T,T^{\prime}$ defined in (1.2)–(1.3), Propositions 1.6 and 1.7 yield
@@ 2.2 Computing the Gaussian approximation
Without loss of generality, fix $t\in T$ with $t\geqslant 0$. We can rewrite $F_{n}(t)$ from 2.2 and compute its derivative: for independent copies $(G_{i},G_{i}^{\prime})$ of
@@
We prove that $p_{t}$ is a well-defined density in Proposition 4.1, and defer the following computation to Sections A.1 and A.1.
Fact 2.2.
@@
$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ -t^{2}\end{bmatrix}$ (2.5)
$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\asymp\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}1&-t\\ -t&\beta\end{bmatrix}.$
@@
Let $q_{t}$ denote the density of $\sqrt{n}\sum_{i=1}^{n}Y_{i}(t)$. By construction $q_{t}$ has mean 00 and covariance $I_{2}$. Moreover, by the change-of-variables formula, it holds
@@
$\varphi(x)\coloneqq\frac{1}{\sqrt{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
We aim to approximate the Kac-Rice integral 2.1 as follows:
$\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y\approx\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{}\left(\det\Sigma_{t}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y\,\mathrm{d}t.$ (2.8)
The validity of this approximation is deferred to Section 3. In the remainder of this section, we solely focus on computing the right hand side integral.
@@
There exists some $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}$ such that
$\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)$ $\asymp\exp\mathopen{}\mathclose{{}\left(-A_{t}-\Theta\mathopen{}\mathclose{{}\left(\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}}\right)y^{2}}\right),$ (2.9)
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}.$
Proof of Lemma 2.3.
Since $\det\Sigma_{t}\asymp\beta^{-2}e^{-t^{2}}$, we have
$\Omega\coloneqq\Sigma_{t}^{-1}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}\beta&t\\ t&1\end{bmatrix}.$
Now as $t\in T$ and so $t^{2}\ll\beta$, we find
@@
$\asymp\Omega_{11}\mu_{t,1}^{2}-2\Omega_{12}\mu_{t,1}(y-\mu_{t,2})+\Omega_{22}(y-\mu_{t,2})^{2}$
$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}\mathopen{}\mathclose{{}\left[\beta-2t\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)+\mathopen{}\mathclose{{}\left(\frac{y}{\mu_{t,1}}+t}\right)^{2}}\right]$
$\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\mu_{t,1}^{2}(\beta-t^{2})+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}$
$\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}+\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}y^{2}.$
This shows the first statement in 2.9. For the second, we have
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{}\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}\right)\,\mathrm{d}y$ $\asymp e^{-A_{t}}\int_{0}^{\infty}ye^{-\Theta\mathopen{}\mathclose{{}\left(\beta^{1/2}e^{t^{2}/{2}}}\right)y^{2}}\,\mathrm{d}y$
$\asymp\beta^{-\frac{1}{2}}e^{-\frac{t^{2}}{2}}e^{-A_{t}}$
by Gaussian integral computations (see A.1). ∎
@@ 2.3 The Kac-Rice integral over $\varphi$
We compute 2.1 under the approximation 2.8. By 2.9 and 2.5, we have that
@@
for any measurable $S\subset\mathbb{R}$. Assuming validity of the Gaussian approximation (see Section 3), it follows from the Kac-Rice formula that the density of modes at $t\in\mathbb{R}$ is proportional to $\sqrt{\beta}e^{-A_{t}}$. We plot this density in Figure 4 with the same choice of $n$ and $\beta$ as in the empirical distribution. We see that they match on the highlighted interval $T$, but not outside of $T$ where the Gaussian approximation—see Remark 3.3.
We compute 2.10 explicitly for $S=T$ and $S=T^{\prime}$.
@@
Proof of Lemma 2.4.
Recall $A_{t}$ from Lemma 2.3. By 2.10, it suffices to show that
@@
As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by 1.2. Let $t_{s}\coloneqq\sqrt{2\log n-s\log\beta}$. As the integrand is positive, for constants $C,C^{\prime}>0$
$\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{2}}^{t_{5/2}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t$
$\geqslant\mathopen{}\mathclose{{}\left(t_{5/2}-t_{2}}\right)\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt_{5/2}^{2}e^{-\frac{t_{2}^{2}}{2}}}\right)$
$\gtrsim\sqrt{\log\beta}\exp\mathopen{}\mathclose{{}\left(-C^{\prime}\beta^{-\frac{1}{2}}\log n}\right)$
@@
as $n,\beta\to\infty$ with $\log n\asymp\log\beta$. Now if $t\in T^{\prime}=[-t_{3},t_{3}]$, we have
$e^{-\frac{t^{2}}{2}}\geqslant e^{-\frac{t_{3}^{2}}{2}}=\beta^{\frac{3}{2}}n^{-1}.$
@@
$\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim\int_{0}^{t_{3}}\exp\mathopen{}\mathclose{{}\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}\right)\,\mathrm{d}t\leqslant\int_{0}^{t_{3}}e^{-Ct^{2}}\,\mathrm{d}t\lesssim 1.\qed$
@@ 3 Leveraging the Edgeworth expansion
In this section, we show the approximation of $p_{t}$ by $\varphi$ is valid in $T$ by showing
@@
One natural idea is to use some asymptotic series to expand $p_{t}$ around $\varphi$, e.g. the Edgeworth expansion, to argue that $|p_{t}-\varphi|\ll\varphi$ in the sense of the integral over $y$. We discuss the two major obstacles we have to overcome in order to implement this approach.
  • Firstly, all known results on validity of asymptotic series such as Edgeworth expansions treat densities $q_{t}$ and $\varphi$ that are independent of $n$. As $\beta$ grows in $n$, we will need to re-derive these results and carefully track the dependence on $\beta$. This will give extra constraints on $(t,\beta,n)$ for the validity of such an asymptotic series. Fortunately, this will be satisfied precisely when $t\in T$.
  • Secondly, even without the $n$-dependence of $\beta$, expanding $q_{t}$ to the order $\varphi$ plus error in for example [BR10, Theorem 19.2] gives error roughly
[…]
    but then the integral over $y$ in 3.1 fails to converge. Therefore, we will need to go to the next term $\psi$ in the Edgeworth series. We control its integral over $y$ and $t$ in 3.1 manually in Section 3.1. Then, we control the higher order terms in Section 3.2 using the approach motivated above, so our error now converges upon integrating over $y$, i.e. roughly
@@
    $\mathopen{}\mathclose{{}\left|p_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right|(\mathbf{x})\lesssim\frac{1}{1+\|\mathbf{x}\|^{3}}\hskip 28.45274pt\text{ for all }\mathbf{x}\in\mathbb{R}^{2}.$
3.1 Bounding the third order error
Recall $Y$ from 2.6. For a multi-index $\alpha\in\mathbb{Z}_{\geqslant 0}^{2}$, let $\kappa^{\alpha}_{t}$ be the cumulant of $Y$ indexed by $\alpha$, which depends on $t$. Let $H^{\alpha}_{t}$ denote the standard Hermite polynomials of with index $\alpha$. The next term in the Edgeworth series is $n^{-1/2}\psi$ with
@@ 3.1 Bounding the third order error
$\psi(\mathbf{x})\coloneqq\frac{\varphi(\mathbf{x})}{6}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}(\mathbf{x})\quad\text{where}\quad H^{\alpha}\coloneqq(-1)^{|\alpha|}\frac{\partial^{\alpha}\varphi}{\varphi}$ (3.2)
If $|\alpha|=3$, $\kappa_{t}^{\alpha}$ are the third moments of $Y$, which we bound in Section A.1.
Fact 3.1.
Let $\eta_{3}$ be the largest third cumulant of $Y$. Then

### v3

@@ 1.4 Notation
We adopt standard notation from asymptotic analysis: we write $f(x)\ll g(x)$ or $f(x)=o(g(x))$ if $f(x)/g(x)\to 0$ as $x\to\infty$; $f(x)\lesssim g(x)$ or $f(x)=O(g(x))$ if there exists a finite, positive constant $C$ such that $f(x)\leqslant Cg(x)$; and we write $f(x)\asymp g(x)$ or $f(x)=\Theta(g(x))$ if $f(x)\lesssim g(x)$ and $g(x)\lesssim f(x)$. We also write $f(x)\sim g(x)$ if $f(x)/g(x)\to 1$ as $x\to\infty$. Similarly, for vector and matrix-valued functions $\mathbf{f}(x)\lesssim\mathbf{g}(x)$ if $f_{i}(x)\lesssim g_{i}(x)$ for every entry, indexed by $i$. We use the analogous notation for $\mathbf{f}(x)\asymp\mathbf{g}(x)$ and $\mathbf{f}(x)\sim\mathbf{g}(x)$. Note that all asymptotic constants are absolute.
@@ 2.1 The Kac-Rice formula
Theorem 2.1 (Kac-Rice, [azais2009level, pp. 62], [adler2009random, Section 11.1]).
@@
    $\mathbb{P}(\upomega(\eta)>\varepsilon)=O(\eta)$ as $\eta\searrow 0^{+}$ for any $\varepsilon>0$, where $\upomega(\cdot)$ denotes the modulus of continuity222defined, for $f:\mathbb{R}\to\mathbb{R}$, as $\upomega(\eta)=\sup_{t,s\colon|t-s|\leqslant\eta}|f(t)-f(s)|$. of $\Psi^{\prime}(\cdot)$.
@@
The Kac-Rice formula extends to any dimension $d\geqslant 1$, and also on manifolds other than $\mathbb{R}^{d}$—see [adler2009random, Section 11.1]. It is the classical tool for computing the expected number of critical points of random fields, with many recent applications including spin glasses [auffinger2013random, fan2021tap] and landscapes of loss functions arising in machine learning [maillard2020landscape]. While the method applies to general densities, the conditional expectation appears infeasible to compute or estimate beyond the Gaussian case. Moreover, we remark that our reliance on the Kac-Rice formula precludes us from any “with high probability” anlogs of Theorem˜1.2, though we do expect such statements to hold.
For the KDE $\widehat{P}_{n}$ defined in ˜1.1, define the random function $F_{n}:\mathbb{R}\to\mathbb{R}$ by
@@
Then $t\in\mathbb{R}$ is an upcrossing of $F_{n}$ at level 0 if and only if $F_{n}(t)=0$ and $F^{\prime}_{n}(t)>0$. This is equivalent to $\widehat{P}^{\prime}_{n}(t)=0$ and $\widehat{P}_{n}^{\prime\prime}(t)<0$, i.e. $t$ is a mode of $\widehat{P}_{n}$. Thus, the number of modes of $\widehat{P}_{n}$ in $T$ is given by $U_{0}(F_{n},T)$. For $T,T^{\prime}$ defined in (1.2)–(1.3), Propositions˜1.7 and 1.8 are equivalent to
@@ 2.2 Computing the Gaussian approximation
Without loss of generality, fix $t\in T$ with $t\geqslant 0$. We can rewrite $F_{n}(t)$ from ˜2.2 and compute its derivative: for independent copies $(G_{i},G_{i}^{\prime})$ of
@@
We prove that $p_{t}$ is a well-defined density in Proposition˜4.1, and defer the following computation to Sections˜5.2 and 5.2.
Lemma 2.2.
@@
$\mu_{t}$ $\coloneqq\sqrt{n}\begin{bmatrix}\mathbb{E}G(t)\\ \mathbb{E}G^{\prime}(t)\end{bmatrix}\sim n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}t\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}-t^{2}\end{bmatrix}$ (2.5)
$\Sigma_{t}$ $\coloneqq\begin{bmatrix}\operatorname{Var}G(t)&\operatorname{Cov}(G(t),G^{\prime}(t))\\ \operatorname{Cov}(G(t),G^{\prime}(t))&\operatorname{Var}G^{\prime}(t)\end{bmatrix}\sim 2^{-\frac{5}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}\begin{bmatrix}2&-t\\ -t&3\beta\end{bmatrix}$.
@@
Let $q_{t}$ denote the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}(t)$. By construction $q_{t}$ has mean 0 and covariance $I_{2}$. Moreover, by the change-of-variables formula, it holds
@@
$\varphi(x)\coloneqq\frac{1}{{2\pi}}e^{-\frac{\|x\|^{2}}{2}}.$
We aim to approximate the Kac-Rice integral ˜2.1 as follows:
$\int_{T}\int_{0}^{\infty}yp_{t}(0,y)\,\mathrm{d}y{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\,\mathrm{d}t}\approx\int_{T}\int_{0}^{\infty}y\mathopen{}\mathclose{{\left(\det\Sigma_{t}}}\right)^{-\frac{1}{2}}\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\,\mathrm{d}t.$ (2.8)
The validity of this approximation is deferred to Section˜3. In the remainder of this section, we solely focus on computing the right hand side integral.
@@
There exists $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}},\delta_{t}\asymp n^{\frac{1}{2}}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}(1-t^{2}/2)$, and $\alpha_{t}\asymp\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}$ such that
$\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}$ $\sim A_{t}+\alpha_{t}(y-\delta_{t})^{2},$ (2.9)
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y$ $\asymp\alpha_{t}^{-1}e^{-A_{t}}.$
Proof of Lemma˜2.3.
We recall ˜2.5 to compute
$\Omega\coloneqq\Sigma_{t}^{-1}\sim 3^{-1}2^{\frac{3}{2}}\beta^{\frac{1}{2}}e^{\frac{t^{2}}{2}}\begin{bmatrix}3\beta&t\\ t&2\end{bmatrix}$
We let the prefactor to be $\alpha_{t}/2$. Since we kept leading coefficients of entries of $\Omega$ and $\mu_{t}$ up to a global absolute constant that we absorb in $\alpha_{t}$ and $\delta_{t}$, it is safe to verify leading coefficients do not cancel and compute asymptotically:
@@
$=\Omega_{11}\mu_{t,1}^{2}-2\Omega_{12}\mu_{t,1}(y-\mu_{t,2})+\Omega_{22}(y-\mu_{t,2})^{2}$
$\sim\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left[\frac{3\beta}{2}-t\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}}}\right)+\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}}}\right)^{2}}}\right]$
$\sim\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left(\frac{3\beta}{2}-\frac{t^{2}}{4}}}\right)+\alpha_{t}\mu_{t,1}^{2}\mathopen{}\mathclose{{\left(\frac{y}{\mu_{t,1}}+\frac{t^{2}-1}{t}-\frac{t}{2}}}\right)^{2}$
$\sim\frac{3}{2}\beta\alpha_{t}\mu_{t,1}^{2}+\alpha_{t}\mathopen{}\mathclose{{\left(y-\frac{\mu_{t,1}}{t}\mathopen{}\mathclose{{\left(1-\frac{t^{2}}{2}}}\right)}}\right)^{2}$
Now, we let $A_{t}$ be the first term and let $\delta_{t}$ be the term subtracting $y$. Verifying the asymptotics of both, we obtain the first statement in ˜2.9. For the second statement, we have
$\int_{0}^{\infty}y\varphi\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y\sim e^{-A_{t}}\int_{0}^{\infty}ye^{-2\alpha_{t}(y-\delta_{t})^{2}}\,\mathrm{d}y\asymp\alpha_{t}^{-1}e^{-A_{t}}$
by a standard fact (Lemma˜5.1) in Gaussian integrals, upon checking $\alpha_{t}^{-1}\delta_{t}^{2}\ll 1$ in our parameter regime of $t\in T$ and $n\lesssim\beta^{2-c}$. The statement and proof of the fact is in Section˜5.1. ∎
@@ 2.3 The Kac-Rice integral over $\varphi$
We compute ˜2.1 under the approximation ˜2.8. By ˜2.9 and ˜2.5, we have that
@@
for any measurable $S\subset\mathbb{R}$. Assuming validity of the Gaussian approximation (see Section˜3), it follows from the Kac-Rice formula that the density of modes at $t\in\mathbb{R}$ is proportional to $\sqrt{\beta}e^{-A_{t}}$. We plot this density in Figure˜4 with the same choice of $n$ and $\beta$ as in the empirical distribution. We see that they match on the highlighted interval $T$, but not outside of $T$ where the Gaussian approximation is no longer valid—see Remark˜3.4.
We compute ˜2.10 explicitly for $S=T$ and $S=T^{\prime}$.
@@
Proof of Lemma˜2.4.
Recall $A_{t}$ from Lemma˜2.3. By ˜2.10, it suffices to show that
@@
As $A_{t}>0$, the integral is at most the length of $T$, which is $O(\sqrt{\log\beta})$ by ˜1.2. Recall that $n\lesssim\beta^{2-c}$. For this constant $c>0$ and $k=1,2$, define
$t_{k}\coloneqq\sqrt{2\log n-\mathopen{}\mathclose{{\left(1+\frac{c}{10^{k}}}}\right)\log\beta}$
Then, $t_{k}>0$ and $t_{k}\in T$ for both $k$. Now, as the integrand is positive, for constants $C,C^{\prime}>0$
$\int_{T}e^{-A_{t}}\,\mathrm{d}t$ $\geqslant\int_{t_{1}}^{t_{2}}\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}}\right)\,\mathrm{d}t$
$\geqslant\mathopen{}\mathclose{{\left(t_{2}-t_{1}}}\right)\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt_{2}^{2}e^{-\frac{t_{1}^{2}}{2}}}}\right)$
$\gtrsim\sqrt{\log\beta}\exp\mathopen{}\mathclose{{\left(-C^{\prime}\beta^{-\frac{1}{2}+\frac{c}{10}}\log n}}\right)$
@@
as $n,\beta\to\infty$ with $\log n\asymp\log\beta$, and $c\in(0,2]$. Now if $t\in T^{\prime}$, we have
$e^{-\frac{t^{2}}{2}}\geqslant\exp\mathopen{}\mathclose{{\left(\log n-\frac{3}{2}\log\beta}}\right)=\beta^{\frac{3}{2}}n^{-1}.$
@@
$\int_{T^{\prime}}e^{-A_{t}}\,\mathrm{d}t\lesssim\int_{T^{\prime}}\exp\mathopen{}\mathclose{{\left(-C\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}}}\right)\,\mathrm{d}t\leqslant\int_{-\infty}^{\infty}e^{-Ct^{2}}\,\mathrm{d}t\lesssim 1.\qed$
@@ 3 Leveraging the Edgeworth expansion
In this section, we show the approximation of $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}$} by $\varphi$ is valid in $T$ by showing
@@
One natural idea is to use some asymptotic series to expand $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}q_{t}$} around $\varphi$, e.g. the Edgeworth expansion, to argue that $|q_{t}-\varphi|\ll\varphi$ in the sense of the integral over $y$. To this end, we cite a standard result in normal approximation theory in the case of identity covariance matrix, that we will follow closely,
Theorem 3.1 ([bhattacharya2010normal, Theorem 19.2 and 19.3]).
Let $X_{n}$ be a sequence of i.i.d. random vectors in $\mathbb{R}^{k}$ with mean zero and identity covariance matrix. Suppose $\mathbb{E}\|X_{1}\|^{s+1}<\infty$ for some integer $s\geqslant 2$, then under suitable conditions, the density $q_{n}$ of $n^{-1/2}(X_{1}+\dots+X_{n})$ admits the asymptotic expansion
@@
$\sup_{\mathbf{x}\in\mathbb{R}^{k}}\mathopen{}\mathclose{{\left(1+\|\mathbf{x}\|^{s}}}\right)\mathopen{}\mathclose{{\left|q_{n}(\mathbf{x})-\sum_{j=0}^{s-2}n^{-\frac{j}{2}}Q_{j}(\mathbf{x})}}\right|\lesssim n^{-\frac{s-1}{2}}$
as $n\to\infty$, where $Q_{j}$ is the $j$-th term of the Edgeworth expansion. In particular, $Q_{0}=\varphi$.
It is tempting to directly apply this theorem with $s=2$ to control $|q_{t}-\varphi|$ in ˜3.1. We discuss the two major obstacles we have to overcome in order to implement this approach.
  • First, Theorem˜3.1 and similar results on validity of asymptotic series such as Edgeworth expansions treat densities $q_{t}$ and $\varphi$ that are independent of $n$. As $\beta$ grows in $n$, we will need to re-derive these results and carefully track the dependence on $\beta$. This will give extra constraints on $(t,\beta,n)$ for the validity of such an asymptotic series. Fortunately, this will be satisfied precisely when $t\in T$. The analog of Theorem˜3.1 in our setting for $s=2,3$ is given as Lemma˜3.5.
  • Second, if we directly apply Theorem˜3.1 with $s=2$ to control $|q_{t}-\varphi|$, then the inner integral over $y\in[0,\infty)$ in ˜3.1 fails to converge as it is of the form
@@
    $\int_{0}^{\infty}\frac{\tilde{y}}{1+\tilde{y}^{2}}\,\mathrm{d}\tilde{y}\quad\text{\quad}\text{where}\quad\tilde{y}\coloneqq\alpha_{t}^{\frac{1}{2}}(y-\delta_{t})$ (3.2)
To overcome this obstacle, we use the above definition of $\tilde{y}$ and Lemma˜2.3 to obtain that
$\mathopen{}\mathclose{{\left\|\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right\|^{2}\sim A_{t}+\tilde{y}^{2}$
This naturally suggests casework on which term has the dominant contribution: for $|\tilde{y}|\leqslant\sqrt{A_{t}}$, we follow Theorem˜3.1 for the $s=3$ case to bound the error of $q_{t}-\varphi$, whereby ˜3.2 integrated from $y=0$ up to $\tilde{y}=\sqrt{A_{t}}$ converges; for $\tilde{y}\geqslant\sqrt{A_{t}}$, we go to the next term $n^{-1/2}\psi$ in the Edgeworth series, and manually control this third order error in Section˜3.1. Finally, we control the higher order terms following Theorem˜3.1 for the $s=3$ case: there, the analog of ˜3.2 for controlling
$\int_{0}^{\infty}y\mathopen{}\mathclose{{\left|q_{t}-\varphi-n^{-\frac{1}{2}}\psi}}\right|\mathopen{}\mathclose{{\left(\Sigma_{t}^{-\frac{1}{2}}[(0,y)-\mu_{t}]}}\right)\,\mathrm{d}y$
converges as $s=3$. This is done in Section˜3.2.
3.1 Bounding the third order error for small $y$
Recall $Y$ from ˜2.6. Let $\varphi$ denote the density of $N(0,I_{2})$ and
$H^{\alpha}(x)\coloneqq(-1)^{|\alpha|}\,\varphi(x)^{-1}\partial^{\alpha}\varphi(x)$
the standard multivariate Hermite polynomials for a multi-index $\alpha\in\mathbb{Z}_{\geqslant 0}^{2}$. Writing $q_{t}$ for the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}(t)$, the third‑order Edgeworth expansion is the multivariate Hermite expansion of $q_{t}/\varphi$:
$\frac{q_{t}(\mathbf{x})}{\varphi(\mathbf{x})}=1+\frac{1}{\sqrt{n}}\sum_{|\alpha|=3}\frac{\kappa_{t}^{\alpha}}{\alpha!}H^{\alpha}(\mathbf{x})+r_{n,t}(\mathbf{x}),$
so the next term is $n^{-1/2}\psi$ with
$\psi(\mathbf{x})=\varphi(\mathbf{x})\sum_{k=0}^{3}\frac{\kappa_{t}^{(k,3-k)}}{k!(3-k)!}H^{(k,\,3-k)}(\mathbf{x}).$ (3.3)
Here $\kappa_{t}^{\alpha}$ denotes the (order-$|\alpha|$) cumulant of the single‑sample vector $Y(t)$, i.e. the $\alpha$‑th mixed derivative at 0 of the cumulant generating function $\log\mathbb{E}e^{\langle u,Y(t)\rangle}$. For $|\alpha|=3$ and our normalization, one has the equivalent moment identity